======20250111====== =====Time===== 1600-0200 =====Attendance===== Dimi, Krishna, Percy, Nick =====Spaces Occupied===== 401, 433c, dome =====Goals===== variable stars =====Equipment Used===== ^ Scope ^ Camera ^ Focus ^ |Main|ASI 2600|| |Finder|ASI 482|2465| =====Frame Table===== ^ Time Observed ^ Source Name ^ Filter ^ Int Time [s] ^ Gain ^ Airmass ^ Mag (filter) ^ RA [h] ^ DEC [deg] ^ Source Type ^Scope ^ Purpose ^ Notes^ | 20250111T1753 | HD_213306 | V | 750ms | 150 | airmass | 4.07 |22 | 58 | type | main | variable star | tracking? Others?| | 20250111T1802 | HIP_117690 | V | 60s | 150 | airmass | 9.90 |23 | 58 | type | main | variable star | tracking? Others?| | 20250111T1827 | HD_236349 | V | 40s | 150 | airmass | 9.15 |16 | 56 | type | main | variable star | tracking? Others?| | 20250111T1840 | HIP_8312 | V | 50s | 150 | airmass | 10.80 | 01 |61 | type | main | variable star | tracking? Others?| | 20250111T1855 | HD_213306 | V | 750ms | 150 | airmass | 4.05 |22 | 58 | type | main | variable star | tracking? Others?| | 20250111T1858 | HIP_117690 | V | 60s | 150 | airmass | 9.90 |23 | 58 | type | main | variable star | tracking? Others?| | 20250111T1912 | HD_236349 | V | 40s | 150 | airmass | 9.15 |15.47 | 56 | type | main | variable star | tracking? Others?| | 20250111T1922 | HIP_8312 | V | 50s | 150 | airmass | 10.80 | 01 |61 | type | main | variable star | tracking? Others?| | 20250111T1937 | HD_278195 | V | 50s | 150 | airmass | 9.75 |05 | 42 | type | main | variable star | tracking? Others?| | 20250111T1957 | HD_31913 | V | 10s | 150 | airmass | 7.5 |05 | 39 | type | main | variable star | tracking? Others?| | 20250111T2001 | HD_213306 | V | 750ms| 150 | airmass | 4.07 |22 | 58 | type | main | variable star | tracking? Others?| | 20250111T2005 | HIP_117690 | V | 60s | 150 | airmass | 9.90 |23 | 58 | type | main | variable star | tracking? Others?| | 20250111T2017 | HD_236349 | V | 40s | 150 | airmass | 9.15 |15.47 | 56 | type | main | variable star | tracking? Others?| | 20250111T2026 | HIP_8312 | V | 50s | 150 | airmass | 10.80 | 01 |61 | type | main | variable star | tracking? Others?| | 20250111T2040 | HD_278195 | V | 50s | 150 | airmass | 9.75 |05 | 42 | type | main | variable star | tracking? Others?| | 20250111T2050 | HD_31913 | V | 10s | 150 | airmass | 7.5 |05 | 39 | type | main | variable star | tracking? Others?| | 20250111T2054 | HD_213306 | V | 750ms | 150 | airmass | 4.07 |22 | 58 | type | main | variable star | tracking? Others?| | 20250111T2058 | HIP_117690 | V | 60s | 150 | airmass | 9.90 |23 | 58 | type | main | variable star | tracking? Others?| | 20250111T2117 | HD_236349 | V | 40s | 150 | airmass | 9.15 |15.47 | 56 | type | main | variable star | tracking? Others?| | 20250111T2136 | HIP_8312 | V | 50s | 150 | airmass | 10.80 | 01 |61 | type | main | variable star | tracking? Others?| | 20250111T2330 | FLATS | n/a | 35s | 150 | airmass | n/a | n/a | n/a | type | main | calibration frame | | 20250111T2322 | DARKS | n/a | 750ms | 150 | airmass | n/a | n/a | n/a | type | main | calibration frame | | 20250111T2323 | DARKS | n/a | 10s | 150 | airmass | n/a | n/a | n/a | type | main | calibration frame | | 20250111T2325 | DARKS | n/a | 35s | 150 | airmass | n/a | n/a | n/a | type | main | calibration frame | | 20250111T2330 | DARKS | n/a | 40s | 150 | airmass | n/a | n/a | n/a | type | main | calibration frame | =====Log===== - Took multiple attempts to open the mirror doors - Pointing to Deneb and Saturn - Slewed and captured HD_213306, HIP_117690, HD_236349, HIP_8312 - Recaptured the above objects again - Slewed and captured HIP_278195, HD_31913 - Recaptured HD_213306, HIP_117690, HD_236349, HIP_8312, HD_278195, HD_31913 - Recaptured HD_213306, HIP_117690 2100 - Percy arrived 2117 - 4th capture of HD_236349 2136 - 4th capture of HIP_8312 * Noticed clouds rolling in mid-capture, around 2140 * This capture is unusable due to cloud interference 2200 - Nick arrived * Received the thumbs up from Brayden to wait the clouds out until 2230 2230 - Still very cloudy, closing the dome to begin taking flat frames 2243 - Flat frames, 35s 2255 - Closed the dome up for darks 2300 - North and south mirror doors stuck, south mirror door went down after tilting the telescope north * waiting until 2310 to see if the north mirror door will close to gravity as well - otherwise we will have to * de-thermalize the dome to thaw the motors 2310 - de-thermalized the dome and the north mirror door went down after 10 mins of thawing 2322 - taking darks =====ToDo===== =====Notes===== =====Weather===== =====Data=====